At the 10% level, do you have enough evidence for the socom 2. A credit card company determined that their current users charge an average of $2500 per year. The company wondered whether giving frequent flyer miles for every purchase would increase card usage. They gave free miles to a sample of 49 credit card customers and found the sample mean to be $2542 with a standard deviation $147 At a- 1%, do you have enough evidence to conclude that frequency flier miles will increase average usage
Answer #1

H0: Null Hypothesis: mu = 2500

HA: Alternative Hypothesis: mu > 2500

SE = s/sqrt{n}

= 147/V49 = 21

Test statistic is:

t = (2542 - 2500)/21 = 2

alpha = 0.01

ndf = 49 - 1 = 48

From Table, critical value of t = 2.4066

Since the calculated value of t = 2 is less than critical value of t = 2.4066, the difference is not significant. Fil to reject null hypothesis.

Conclusion:

The data do not support the claim that the frequency flier miles will increase average usage.